OpenStudy (anonymous):

what is the equation of the line with slope of -2 that passes through (0,5)? please explained.

OpenStudy (anonymous):

use the slope-point formula

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

plug in your slope into m and point into (x1,y1)

OpenStudy (anonymous):

The equation of the point-slope form is as follows: \[y _{1} - y = m (x _{1} -x)\] Plug in (0,5) for the x and y values.

OpenStudy (anonymous):

i still don't understand.

OpenStudy (anonymous):

i came up with y=-2x+5?

OpenStudy (anonymous):

yep

hero (hero):

y = mx + b 5 = -2(0) + b 5 = b y = -2x + 5